OpenStudy (anonymous):

what is the value of tan22.5 + cot22.5 = ? without using calculator !

OpenStudy (anonymous):

a)sqrt(3) /2 b)2 c)sqrt(2) d) 2sqrt(2)

OpenStudy (shamim):

\[\tan 22.5+\cot 22.5\]

OpenStudy (shamim):

\[\tan 22.5+\frac{ 1 }{ \tan 22.5 }\]

OpenStudy (shamim):

r u following me

OpenStudy (anonymous):

ok i got 2 cosec2 times 22.5

OpenStudy (shamim):

\[\frac{ \tan ^{2}22.5 +1}{ \tan 22.5 }\]

OpenStudy (anonymous):

i got the answer it's 2 sqrt 2 isn't it !

OpenStudy (shamim):

so wts ur problem

OpenStudy (anonymous):

thanks @shamim

OpenStudy (shamim):

i think u got result

OpenStudy (shamim):

is not it

OpenStudy (anonymous):

i got lil bit confused !

OpenStudy (shamim):

why

OpenStudy (shamim):

u got \[\frac{ 2 }{ \sin45 }\]

OpenStudy (shamim):

is not it

OpenStudy (shamim):

\[\frac{ 2 }{ \frac{ 1 }{ \sqrt{2} } }\]

OpenStudy (shamim):

\[2\sqrt{2}\]

OpenStudy (anonymous):

yeh i got the same !

OpenStudy (anonymous):

can you tell me how you did it ! @shamim

OpenStudy (shamim):

ya

OpenStudy (shamim):

ya i can tell u

OpenStudy (shamim):

\[\frac{ \sec ^{2}22.5 }{ \tan 22.5 }\]

OpenStudy (anonymous):

yes then ?

OpenStudy (shamim):

\[\frac{ \frac{ 1 }{ \cos ^{2}22.5 } }{ \frac{ \sin 22.5 }{ \cos 22.5 } }\]

OpenStudy (shamim):

\[\frac{ 1 }{ \cos ^{2}22.5 } \times \frac{ \cos 22.5 }{ \sin 22.5 }\]

OpenStudy (shamim):

\[\frac{ 1 }{ \cos 22.5 \sin 22.5 }\]

OpenStudy (shamim):

\[\frac{ 2 }{ 2 \sin 22.5 \cos 22.5 }\]

OpenStudy (shamim):

\[\frac{ 2 }{ \sin 2 \times 22.5 }=\frac{ 2 }{ \sin 45 }\]

OpenStudy (anonymous):

thank you very much !

OpenStudy (shamim):

u r most welcome